mathslover (mathslover):

anti derivative of x^5 + x^4

OpenStudy (lgbasallote):

use power formula \[\huge \int x^n dx \implies \frac{x^{n+1}}{n+1}\] does that help?

mathslover (mathslover):

\[\large{d/dx ( x^5+ x^4)}\] \[\large{d/dx(x^4(x+1))}\] \[\large{d/dx(x^4)*d/dx(x+1)}\] \[\large{4x^3 *2 = 8x^3}\]

mathslover (mathslover):

ok so i have to do : \[\int{x^5+x^4}\]

OpenStudy (anonymous):

anti derivative means integration

OpenStudy (anonymous):

ya

mathslover (mathslover):